Find any relative maxima or minima for f(x)=x^2 +2x-8
Answer by Earlsdon(6294)

The function: f(x) = x^2 + 2x - 8 is that of a parabola that opens upward (the coefficient of x^2 is positive (+1))and therefore has a vertex that is a minimum.
The x-coordinate of the vertex is given by: x = -b/2a
So the x-coordinate of the vertex is: x = -(2)/2(1) = -1 Substitute this into the equation to find the y-coordinte: y = (-1)^2 + 2(-1) - 8 = 1 - 2 - 8 = 9
So the minimum (vertex) is at (-1, 9)